OpenStudy (anonymous):

Lim (x^2/9 - 3/x) x->27 i plugged 27 in as x and it says it's wrong and to keep 27 as (9)(3)

zepdrix (zepdrix):

This is the expression?\[\Large\bf\sf \lim_{x\to27}\frac{x^2}{9}-\frac{3}{x}\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Oh that would have been easier actually... the hint that they used.

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to9\cdot3}\frac{x^2}{9}-\frac{3}{x}\quad=\quad \frac{(9\cdot3)^2}{9}-\frac{3}{9\cdot3}\]

zepdrix (zepdrix):

\[\Large\bf\sf =\quad \frac{9^2\cdot3^2}{9}-\frac{3}{9\cdot3}\]

zepdrix (zepdrix):

Any confusion on that? :o

zepdrix (zepdrix):

Understand how to simplify from there?

OpenStudy (anonymous):

i put the answer as 80.88 after simplifying and it's wrong

zepdrix (zepdrix):

Don't approximate. Decimals are bad bad bad. Keep the final answer as a fraction. Or do they prefer a decimal for your online answer?

OpenStudy (anonymous):

oh i just saw simplified fraction

OpenStudy (anonymous):

so should i just put the answer as 81-1/9 sorry i only have one try left before i get it wrong

zepdrix (zepdrix):

One try? Let's not 81-1/9 then. We don't wanna chance it.

zepdrix (zepdrix):

So we need a common denominator.\[\Large\bf\sf \frac{81\cdot 9}{9}-\frac{1}{9}\]

zepdrix (zepdrix):

What do you get when you simplify that down? :x

OpenStudy (anonymous):

728/9

zepdrix (zepdrix):

Yayyy good job \c:/

OpenStudy (anonymous):

thank you